# Trivariate Bicycle Codes

### v1

@@ Appendix C Tanner Graph Details
Before diving into any technical details, it is important to point out that the physical implementation of a QECC does not have to directly conform to the simplest Tanner Graph structure, as along as the graph connectivity is retained. One can arbitrarily permute the qubits around physical space as they wish, changing the physical implementation of the code and the corresponding stabilizers’ supports. However, a symmetric Tanner Graph structure can greatly ease hardware implementation, by noting that the Tanner Graph embedded in D-dimensional Euclidean Space $\mathbb{R}^{D}$ directly suggests one choice of physical implementation, formalized by Proposition 5. Therefore, we seek the simplest possible embedding in physical space, in our case the toric layout.
@@
Given a graph G, we now rewrite vertices of G into indices $\{i\}$, and the edges of G into a matrix, called the adjacency matrix A, where $a_{ij}=1$ when there is an edge between vertex i and vertex j, 0 otherwise. Thus, the $A_{i}$ and $B_{j}$ terms that make up A and B in our definition of a BB-QLDPC code, give rise to the adjacency matrix of the code’s Tanner Graph.
@@
Define a qubit two-block CSS code as CSS codes such that $H_{X}=[A|B]$ and $H_{Z}=[B^{T}|A^{T}]$, where A and B commute and have the same size. Two-blocks allow us to partition data qubits into L and R qubits. BB codes are a special type of two-block CSS code, amongst others that have been studied [58].
@@
We choose the physical embedding E of qubits provided in the hypothesis, such that each stabilizer S is supported in its ball of radius $r_{S}$, specifically on those data qubit vertices that share an edge with the check vertex of S. In our paper, this refers to the toric layout embedded on a lattice in $S^{1}\times S^{1}$, and indeed all for our weight-5 [[30,4,5]] code in Fig 1, our stabilizers will be supported on 4 neighbouring qubit sites, and 1 long-range qubit site. ∎
Note that the above proposition also implies that for any QECC such that its Tanner Graph has a local embedding in $\mathbb{R}^{D}$, the QECC parameters [[n,k,d]] must satisfy the Bravyi-Poulin-Terhal bound [45], the statement that for geometrically local codes, $\exists$ geometric locality constant $c$ such that $kd^{\frac{2}{D-1}}\leq cn$. This bound holds even if one chooses a non-local physical implementation. Thus, one should remain cautious about the performance of a code, even when provided with a non-local embedding of its Tanner Graph, as it may be graph isomorphic to a geometrically local code, where the notion of a graph isomorphism is defined below in Definition 6.
@@
This allows a bi-partition into two degree-3 subgraphs, allowing us to adapt the proof of [20] on each degree-3 subgraph of our weight-6 codes mutatis mutandis. ∎
Note that the above proof breaks down if we have A = $A_{1}+A_{3}$, B = $B_{1}+B_{2}+B_{3}+B_{4}+B_{5}$ such as in our weight-7 code. It is easy to extend the above argument to show that weight-7,8,9 codes have a thickness-3 implementation (and analogously for larger weights and thickness), by splitting into multiple CSS subcodes, but high thickness is not desirable for near-term hardware implementation. The Tanner Graph is degree-7, but it is known that degree-7 thickness-2 graphs exist, such as the complete $K_{7}$ graph, defined as having 7 vertices and all-to-all connectivity [42]. Furthermore, a lower bound on graph thickness is obtained from a corollary of Euler’s polyhedron formula, that given a graph with $p$ vertices and $q$ edges, $\theta\geq\frac{q}{3p-6}$ [59], yielding a lower bound of $\theta\geq\frac{25}{29}$ for our weight-7 [[30, 4, 5]] codes. It is easy to verify that the graph is not planar, but we could remain hopeful that it has thickness $\theta=2$.
In order to rule out the possibility that the graph may fortunately be of thickness $\theta\!=\!2$, a numerical program is necessary. Furthermore, by explicitly checking subgraphs, we will obtain an bi-planar layout for experimental realization. This brute-force procedure for checking $\theta\!=\!2$ is generically NP-complete [60]. For very small codes such as the Shor Code [6], it is still practical to enumerate decompositions into 2 subgraphs, and then check the planarity of each subgraph. However, our weight-7 [[30, 4, 5]] code did not provide an answer after a few days of simulation. An upper bound provided below on the computational complexity of a brute-force search for bi-planarity suggests that the difficulty of iterating over subgraphs is the key area for improvement in aiming towards a $\text{poly}(n)$ time complexity.
@@
Intuitively, we can relabel the vertices and edges of G to get those of H. This is what allows us to view the Tanner Graph of TB-QLDPC codes on the discrete torus.
@@
We first define two concepts relevant in our BB codes.
@@
Recall that the BB codes are defined using shift matrices x, y, z = $S_{l}\otimes I_{m}$, $I_{l}\otimes S_{m}$, $S_{l}\otimes S_{m}$. $S_{l}$ and $S_{m}$ are circulant matrices. Thus x, y, z are cyclic matrices.
@@
By this definition, the codes with a toric layout such as in Fig 1 will have $t_{x}$ = $t_{y}$ = 0, the trivial twist. Non-trivial twisted toric layouts no longer have global translational symmetry, but retain it in the bulk. Some work has been done on studying twisted layouts [43], and it can be fruitful to find a systematic way to incorporate such theoretical results in a search for desirable weight-4 QLDPC codes for near-term implementation.
@@
A QLDPC code C has a ($\sigma$, $\tau$)-tangled toric layout $\iff\exists\sigma\in S_{2\mu},\tau\in S_{2\lambda}$ where $S_{q}$ is the permutation group on $q$ elements, such that the Tanner Graph of C has a toric layout after applying $\sigma$ along the x-direction of the periodic boundary conditions of the toric layout, similarly for $\tau$ along the y-direction.
@@
We encourage readers to find systematic ways of studying tangling parameters $(\sigma,\tau)$ of codes that are surface codes in the bulk. As a first step in this direction, we provide one algorithm below for weight-4 BB Q-LDPC codes that have a tangled toric layout.
@@
For any two-block CSS Q-LDPC code of sparsity 4, suppose it has a tangled toric layout known to have torus parameters $\mu$,$\lambda$.

### v2

@@ Appendix F Tanner Graph Details
Before diving into any technical details, it is important to point out that the physical implementation of a QECC does not have to directly conform to the simplest Tanner Graph structure, as along as the graph connectivity is retained. One can arbitrarily permute the qubits around physical space as they wish, changing the physical implementation of the code and the corresponding stabilizers’ supports. However, a symmetric Tanner Graph structure can greatly ease hardware implementation, by noting that the Tanner Graph embedded in D-dimensional Euclidean Space $\mathbb{R}^{D}$ directly suggests one choice of physical implementation, formalized by Proposition 5. Therefore, we seek the simplest possible embedding in physical space, which in the case of MB codes is the toric layout.
@@
Given a graph G, we now rewrite vertices of G into indices $\{i\}$, and the edges of G into a matrix, called the adjacency matrix A, where $a_{ij}=1$ when there is an edge between vertex i and vertex j, 0 otherwise. Thus, the $A_{i}$ and $B_{j}$ terms that make up A and B in our definition of a MB-QLDPC code, give rise to the adjacency matrix of the code’s Tanner Graph.
@@
Define a qubit two-block CSS code as CSS codes such that $H_{X}=[A|B]$ and $H_{Z}=[B^{T}|A^{T}]$, where A and B commute and have the same size. Two-blocks allow us to partition data qubits into L and R qubits. TB codes are a special type of two-block CSS code, amongst others that have been studied where A and B arise from a group algebra [58].
@@
We choose the physical embedding E of qubits provided in the hypothesis, such that each stabilizer S is supported in its ball of radius $r_{S}$, specifically on those data qubit vertices that share an edge with the check vertex of S. In our paper, this refers to the toric layout embedded on a lattice in $S^{1}\times S^{1}$, and indeed all for our weight-5 $\llbracket 30,4,5\rrbracket$ code in Fig 1, our stabilizers will be supported on 4 neighbouring qubit sites, and 1 long-range qubit site. ∎
Note that the above proposition also implies that for any QECC such that its Tanner Graph has a local embedding in $\mathbb{R}^{D}$, the QECC parameters $\llbracket n,k,d\rrbracket$ must satisfy the Bravyi-Poulin-Terhal bound [27], the statement that for geometrically local codes, $\exists$ geometric locality constant $c$ such that $kd^{\frac{2}{D-1}}\leq cn$. This bound holds even if one chooses a non-local physical implementation. Thus, one should remain cautious about the performance of a code, even when provided with a non-local embedding of its Tanner Graph, as it may be graph isomorphic to a geometrically local code, where the notion of a graph isomorphism is defined below in Definition 5.
@@
This allows a bi-partition into two degree-3 subgraphs, allowing us to adapt the proof of [20] on each degree-3 subgraph of our weight-6 codes. ∎
Note that the above proof breaks down if we have $A=A_{1}+A_{3},B=B_{1}+B_{2}+B_{3}+B_{4}+B_{5}$ such as in our weight-7 code. It is easy to extend the above argument to show that weight-7,8,9 codes have a thickness-3 implementation (and analogously for larger weights and thickness), by splitting into 3 CSS subcodes, but high thickness is not desirable for near-term hardware implementation. The Tanner Graph is degree-7, but it is known that degree-7 thickness-2 graphs exist, such as the complete $K_{7}$ graph, defined as having 7 vertices and all-to-all connectivity [57]. Furthermore, a lower bound on graph thickness is obtained from a corollary of Euler’s polyhedron formula, that given a graph with $p$ vertices and $q$ edges, $\theta\geq\frac{q}{3p-6}$ [59], yielding a lower bound of $\theta\geq\frac{25}{29}$ for our weight-7 $\llbracket 30,4,5\rrbracket$ codes. It is easy to verify that the graph is not planar, but we could remain hopeful that it has thickness $\theta=2$.
In order to rule out the possibility that the graph may fortunately be of thickness $\theta\!=\!2$, a numerical program is necessary. Furthermore, by explicitly checking subgraphs, we will obtain an bi-planar layout for experimental realization. This brute-force procedure for checking $\theta\!=\!2$ is generically NP-complete [60]. For very small codes such as the Shor Code [6], it is still practical to enumerate decompositions into 2 subgraphs, and then check the planarity of each subgraph. However, our weight-7 $\llbracket 30,4,5\rrbracket$ code did not provide an answer after a few days of simulation. An upper bound provided below on the computational complexity of a brute-force search for bi-planarity suggests that the difficulty of iterating over subgraphs is the key area for improvement in aiming towards a $\text{poly}(n)$ time complexity.
@@
At this point, it is worth noting that we could generically define a $b$-block group algebra code as $[A_{1}|A_{2}|...|A_{b}]$, enforce that $\sum_{i}A_{i}A_{k-i+1}^{T}=0$ to enforce the CSS condition, and study it as a qLDPC code. We expect the generic framework to carry over.
@@
Intuitively, we can relabel the vertices and edges of G to get those of H. This is what allows us to view the Tanner Graph of MB-QLDPC codes on the discrete torus.
@@
  1. 1.
    $\langle A_{i}A_{j}^{T},B_{g}B_{h}^{T}\rangle=M$ and
  2. 2.
    $\operatorname{ord}(A_{i}A_{j}^{T})\operatorname{ord}(B_{g}B_{h}^{T})=lm$.
To explain the idea behind Proposition 8, which gives our criterion for a toric layout, split our data qubits into two blocks through the definition below.
$H_{X}=[A|B]\quad\text{and}\quad H_{Z}=[B^{T}|A^{T}]$
L represents first half of the qubits (left side of block matrix), R represents the second half (right side of block matrix).
[…]
@@
We first define two concepts relevant in our TB codes.
@@
Recall that the TB codes are defined using shift matrices x, y, z = $S_{l}\otimes I_{m}$, $I_{l}\otimes S_{m}$, $S_{l}\otimes S_{m}$. $S_{l}$ and $S_{m}$ are circulant matrices. Thus x, y, z are cyclic matrices.
@@
By this definition, the codes with a toric layout such as in Fig 1 will have $t_{x}$ = $t_{y}$ = 0, the trivial twist. Non-trivial twisted toric layouts no longer have global translational symmetry, but retain it in the bulk. Some work has been done on studying twisted layouts [62], and it can be fruitful to find a systematic way to incorporate such theoretical results in a search for desirable weight-4 QLDPC codes for near-term implementation, as a first step beyond the toric code.
@@
A CSS QLDPC code C has a ($\sigma$, $\tau$)-tangled toric layout $\iff\exists\sigma\in S_{2\mu},\tau\in S_{2\lambda}$ where $S_{q}$ is the permutation group on $q$ elements, such that the Tanner Graph of C has a toric layout after applying $\sigma$ along the x-direction of the periodic boundary conditions of the toric layout, similarly for $\tau$ along the y-direction.
@@
We encourage readers to find systematic ways of studying tangling parameters $(\sigma,\tau)$ of codes that are surface codes in the bulk. As a first step in this direction, we provide one algorithm below for weight-4 TB codes that have a tangled toric layout.
@@
For any two-block CSS QLDPC code of sparsity 4 with equal X and Z degree per data qubit (2 each), suppose it has a tangled toric layout known to have torus parameters $\mu$,$\lambda$.